a collection of 36 dimes and quarters is worth 4 dollars and 80 cents; how many dimes are there?....please I would like help over this I have a test tommorow and cant fail it or I am off the baseball team...please help
Answer by scott8148(6628)

"36 dimes and quarters" ___ d + q = 36 ___ 10d + 10q = 360 (A)

"worth 4 dollars and 80 cents" ___ 10d + 25q = 480 (B)

subtracting equation (A) from equation (B) to eliminate "d" ___ 15q = 120 ___ q = 8

substituting ___ d + (8) = 36 ___ d = 28